The Math Pumpkin football team played a game against the Math Wizards. They played two halves, each of which lasted twenty-four minutes. The Math Pumpkin team’s time of possession was four minutes more than the Math Wizard’s. How long was the Math Wizard’s time of possession?
Answer by greenestamps(13195)



(1) Using formal algebra....

x = minutes of possession for Math Wizards
x+4 = minutes of possession for Math Pumpkin

The total time was 48 minutes (two 24-minute halves):
x+x+4=48
2x+4=48
2x=44
x=22

ANSWER: The Math Wizards' time of possession was x=22 minutes

(2) Using logical reasoning and simple mental arithmetic....

If the time of possession had been the same for both teams, each team would have had possession for 24 minutes. For one team to have 4 minutes of possession more than the other team, the times of possession must have been 24+2=26 and 24-2=22 minutes.

ANSWER (again): The Math Wizards' time of possession was x=22 minutes